# Three elliptic closed characteristics on the non-degenerate compact convex hypersurfaces in $\mathbb{R}^{6}$

### 13 Mar 2026 v1

@@ Abstract
Let $\Sigma\subset\mathbb{R}^{2n}$ with $n\geq 2$ be any $C^{2}$ compact convex hypersurface. The stability of closed characteristics has attracted considerable attention in related research fields. A long-standing conjecture states that all closed characteristics are irrationally elliptic, provided $\Sigma$ possesses only finitely geometrically distinct closed characteristics. This conjecture has been fully resolved only in $\mathbb{R}^{4}$, while it remains completely open in higher dimensions. Even in $\mathbb{R}^{6}$, it is unknown whether there exist three elliptic closed characteristics. In this paper, we first prove that for any non-degenerate $\Sigma\in\mathbb{R}^{2n}$ with finitely many closed characteristics, at least two of them are irrationally elliptic. Moreover, for any non-degenerate $\Sigma\subset\mathbb{R}^{6}$ with finitely many closed characteristics, we obtain at least three elliptic characteristics, of which at least two are irrationally elliptic. Based on the $n$-or-$\infty$ conjecture, three elliptic closed characteristics are optimal. This result provide theoretical support for further research on this conjecture.
AMS Subject Classification: 58E05, 37J45, 34C25
@@ 1 Introduction and main results
is called the associated symplectic path of $(\tau,x)$. The eigenvalues of $\gamma_{x}(\tau)$ are called Floquet multipliers, and denote the elliptic height $e(\gamma_{x}(\tau))$ of $\gamma_{x}(\tau)$ to be the total algebraic multiplicity of all eigenvalues of $\gamma_{x}(\tau)$ on the unit circle $\mathbb{U}=\{z\in\mathbb{C}\,|\,|z|=1\}$. As usual, a closed characteristic $(\tau,x)$ is non-degenerate, if $1\in\sigma(x)$ is of algebraic multiplicity $2$; a hypersurface $\Sigma$ is called non-degenerate, if all the closed characteristics on the given hypersurface together with all of their iterations are non-degenerate as periodic solutions of the corresponding Hamiltonian system. A closed characteristic $(\tau,x)$ is called elliptic, if $e(\gamma_{x}(\tau))=2n$; hyperbolic if $1$ is a double Floquet multiplier of it and $e(\gamma_{x}(\tau))=2$; irrationally elliptic if in an appropriate coordinate $\gamma_{x}(\tau)$ is the $\diamond$-product of one $N_{1}(1,1)=\left(\begin{array}[]{cc}1&1\\ 0&1\\ \end{array}\right)$ and n1n-1 rotation matrices $\left(\begin{array}[]{cc}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{array}\right)$ with rotation angles $\theta_{i}$ being irrational multiples of $\pi$ for $1\leq i\leq n-1$. Hence, the stability problem refers to estimating the number of elliptic closed characteristics that may exist for system (1.3).
@@
For any non-degenerate surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then at least two closed characteristics are irrational elliptic.
@@
By combining the methods of [1], [11] and this paper, the result of Theorem 1.2 can perhaps be generalized to dynamically convex Reeb flow on non-degenerate star-shaped surfaces.
@@
How to remove the non-degeneracy condition in Theorem 1.2 is an interesting question, but for now, we are unable to do so. In fact, it is still unknown whether the following Assumption A holds, except for the case $n=2$. We believe it to be true for $n\geq 3$, but we are currently unable to prove it; perhaps a deeper analysis of the methods in [11] might lead to progress.
@@ 2.1 Definition of the Maslov-type index
and denote by $M^{\diamond k}$ the $k$-fold $\diamond$-product $M\diamond\cdots\diamond M$. Note that the $\diamond$-product of any two symplectic matrices is symplectic. For any two paths $\gamma_{j}\in\mathcal{P}_{\tau}(2n_{j})$ with $j=0$ and $1$, let $\gamma_{0}\diamond\gamma_{1}(t)=\gamma_{0}(t)\diamond\gamma_{1}(t)$ for all $t\in[0,\tau]$.
@@ 2.2 Iterative formula of the Maslov-type index
Here we call $N_{2}(\omega,b)$ nontrivial if $(b_{2}-b_{3})\sin\theta<0$ and trivial if $(b_{2}-b_{3})\sin\theta>0$. For any $M_{i}\in\emph{Sp}(2n_{i})$ with $i=0$ and $1$, there holds
@@ 2.3 Common index jump theorem
where $\epsilon\in(0,\min\{\delta,1/3\})$, and $\chi_{k},\chi_{k,j}=0$ or $1$ for $k=1,\dots,q$ and $j=1,\dots,\mu_{1},2,\dots,\mu_{2},\dots,\mu_{q}$. In fact, such a vertex $\chi$ exists, as stated in Theorem 2.2.
Theorem 2.2.
@@
Given $M_{0}\in\mathbb{N}$, by the proof of Theorem 4.1 of [19], we may further require $M_{0}|N$ (since the closure of the set $\{\{Nv\}:N\in\mathbb{N},M_{0}|N\}$) is still a closed additive subgroup of $T^{h}$ for some $h\in\mathbb{N}$. Then we can use the step 2 in Theorem 4.1 of of [19] to get $N$), hence in our choice of $(N,m_{1},...,m_{q})$ in Theorem 2.1, we can choose $M_{0}$ good enough such that $N\in\mathbb{N}$ further satisfies
@@
(cf.[19]) For some $a\in A(v)$, let $(N,m_{1},\cdots,m_{q})\in\mathbb{N}^{q+1}$ be given in Theorem 2.1 and 2.2. Define $\varrho_{n}(\Sigma):\mathcal{H}(2n)\to\mathbb{N}\cup\{+\infty\}$ by
@@ 3 The elliptic closed characteristics
$r\delta<\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\left\{m_{j(2)}\frac{\theta_{j}}{\pi}\right\},$ (3.11)
@@
For any fix $a\in A(v)$ and $s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$, consider the inject map $p(N-s+1)=([(\tau_{j(s)},x_{j(s)})],2m_{j(s)})$ given by Theorem 2.4. Now, from Lemma 3.3, we can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$. Assume that the compact convex surface $\Sigma$ is non-degenerate, then $(1)$ in Theorem 3.6 implies $x_{j(1)}$ is irrationally elliptic. Since $\chi_{j(1)}(\hat{a})=1$, from $(2)$ in Theorem 3.6, we know $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$. Now, choose $\tilde{a}=-\hat{a}\in A(v)$, using Theorem 2.4 again, we still have $(\tilde{N},\tilde{m}_{1},\ldots,\tilde{m}_{q})$ , $\tilde{j}(s)$ and inject map $p(\tilde{N}-s+1)=([(\tau_{\tilde{j}(s)},x_{\tilde{j}(s)})],2\tilde{m}_{\tilde{j}(s)}),s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$. Similar, $(1)$ in Theorem 3.6 implies $x_{\tilde{j}(1)}$ is irrationally elliptic. If $\tilde{j}(1)\neq j(1)$, then $x_{j(1)},x_{\tilde{j}(1)}$ are two different irrationally elliptic closed characteristics. If $\tilde{j}(1)=j(1)$, from the definition of $\chi(a)$, we know that $\chi_{j(1)}(\hat{a})=1$ implies $\chi_{j(1)}(-\hat{a})=0$, hence $\chi_{\tilde{j}(1)}(\tilde{a})=\chi_{j(1)}(-\hat{a})=0$, but $\hat{i}(x_{\tilde{j}(1)},1)=\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$, this contradict with $(2)$ in Theorem 3.6, hence this case will not happen, we always have $\tilde{j}(1)\neq j(1)$ and we get two irrationally elliptic closed characteristics. This completes the proof. ∎
@@
$\gamma_{j(2)}(\tau)$ $\simeq N_{1}(1,1)\diamond R(\theta_{1})\diamond R(\theta_{2}).$
Since $\Sigma$ is non-degenerate, $x_{j(2)}$ must be irrational elliptic. In summary, we obtain that $x_{j(1)}$ is irrational elliptic and either $x_{j(2)}$ or $x_{j(3)}$ must be irrational elliptic. This completes the proof. ∎

### 16 Mar 2026 v2

@@ Abstract
Let $\Sigma\subset\mathbb{R}^{2n}$ with $n\geq 2$ be any $C^{2}$ compact convex hypersurface. The stability of closed characteristics has attracted considerable attention in related research fields. A long-standing conjecture states that all closed characteristics are irrationally elliptic, provided $\Sigma$ possesses only finitely geometrically distinct closed characteristics. This conjecture has been fully resolved only in $\mathbb{R}^{4}$, while it remains completely open in higher dimensions. Even in $\mathbb{R}^{6}$, it is unknown whether there exist three elliptic closed characteristics. In this paper, we first prove that for any $\Sigma\subset\mathbb{R}^{2n}$ with finitely many closed characteristics, there exist at least two elliptic closed characteristics, which possess a nice symplectic normal form. In particular, as a simple corollary, they are irrational elliptic when $\Sigma$ is non-degenerate. Moreover, for any non-degenerate $\Sigma\subset\mathbb{R}^{6}$ with finitely many closed characteristics, we obtain at least three elliptic characteristics, of which at least two are irrationally elliptic. Based on the $n$-or-$\infty$ conjecture, three elliptic closed characteristics are optimal. This result provide theoretical support for further research on this conjecture.
AMS Subject Classification: 58E05, 53D12, 34C25
@@ 1 Introduction and main results
is called the associated symplectic path of $(\tau,x)$. The eigenvalues of $\gamma_{x}(\tau)$ are called Floquet multipliers, and denote the elliptic height $e(\gamma_{x}(\tau))$ of $\gamma_{x}(\tau)$ to be the total algebraic multiplicity of all eigenvalues of $\gamma_{x}(\tau)$ on the unit circle $\mathbb{U}=\{z\in\mathbb{C}\,|\,|z|=1\}$. As usual, a closed characteristic $(\tau,x)$ is non-degenerate, if $1\in\sigma(x)$ is of algebraic multiplicity 22; a hypersurface $\Sigma$ is called non-degenerate, if all the closed characteristics on the given hypersurface together with all of their iterations are non-degenerate as periodic solutions of the corresponding Hamiltonian system. A closed characteristic $(\tau,x)$ is called elliptic, if $e(\gamma_{x}(\tau))=2n$; hyperbolic if 11 is a double Floquet multiplier of it and $e(\gamma_{x}(\tau))=2$; irrationally elliptic if in an appropriate coordinate $\gamma_{x}(\tau)$ is the $\diamond$-product of one $N_{1}(1,1)=\left(\begin{array}[]{cc}1&1\\ 0&1\\ \end{array}\right)$ and n1n-1 rotation matrices $\left(\begin{array}[]{cc}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{array}\right)$ with rotation angles $\theta_{i}$ being irrational multiples of $\pi$ for $1\leq i\leq n-1$. Hence, the stability problem refers to estimating the number of elliptic closed characteristics that may exist for system (1.3).
@@
For any surfaces $\Sigma\in\mathcal{H}(2n)$ with $^{\#}\widetilde{\mathcal{J}}(\Sigma)<+\infty$, then at least two closed characteristics are elliptic whose momodormy matrix possess a nice symplectic normal form
$\gamma(\tau)$ $\simeq N_{1}(1,1)\diamond I_{2p_{0}}\diamond N_{1}(1,-1)^{\diamond p_{+}}\diamond N_{1}(-1,1)^{\diamond q_{-}}\diamond-I_{2q_{0}}$
with all $N_{2}(\lambda_{1},\nu_{1}),\ldots,N_{2}(\lambda_{r_{0}},\nu_{r_{0}})$ are rational normal form. Moreover, when $\Sigma$ is non-degenerate, then they are irrationally elliptic.
@@
By combining the methods of [1], [11] and this paper, the result of Theorem 1.2 can perhaps be generalized to dynamically convex Reeb flow on star-shaped surfaces.
@@
How to obtain the irrationally elliptic without the non-degeneracy condition in Theorem 1.2 is an interesting question, but for now, we are unable to do so. In fact, it is still unknown whether the following Assumption A holds, except for the case $n=2$. We believe it to be true for $n\geq 3$, but we are currently unable to prove it; perhaps a deeper analysis of the methods in [11] might lead to progress.
@@ 2.1 Definition of the Maslov-type index
and denote by $M^{\diamond k}$ the $k$-fold $\diamond$-product $M\diamond\cdots\diamond M$. Note that the $\diamond$-product of any two symplectic matrices is symplectic. For any two paths $\gamma_{j}\in\mathcal{P}_{\tau}(2n_{j})$ with $j=0$ and 11, let $\gamma_{0}\diamond\gamma_{1}(t)=\gamma_{0}(t)\diamond\gamma_{1}(t)$ for all $t\in[0,\tau]$.
@@ 2.2 Iterative formula of the Maslov-type index
Here we call $N_{2}(\omega,b)$ nontrivial if $(b_{2}-b_{3})\sin\theta<0$ and trivial if $(b_{2}-b_{3})\sin\theta>0$. For any $M_{i}\in\emph{Sp}(2n_{i})$ with $i=0$ and 11, there holds
@@ 2.3 Common index jump theorem
where $\epsilon\in(0,\min\{\delta,1/3\})$, and $\chi_{k},\chi_{k,j}=0$ or 11 for $k=1,\dots,q$ and $j=1,\dots,\mu_{1},2,\dots,\mu_{2},\dots,\mu_{q}$. In fact, such a vertex $\chi$ exists, as stated in Theorem $2.$
Theorem $2.$
@@
Given $M_{0}\in\mathbb{N}$, by the proof of Theorem 4.1 of [19], we may further require $M_{0}|N$ (since the closure of the set $\{\{Nv\}:N\in\mathbb{N},M_{0}|N\}$) is still a closed additive subgroup of $T^{h}$ for some $h\in\mathbb{N}$. Then we can use the step 2 in Theorem 4.1 of of [19] to get NN), hence in our choice of $(N,m_{1},...,m_{q})$ in Theorem 2.1, we can choose $M_{0}$ good enough such that $N\in\mathbb{N}$ further satisfies
@@
(cf.[19]) For some $a\in A(v)$, let $(N,m_{1},\cdots,m_{q})\in\mathbb{N}^{q+1}$ be given in Theorem 2.1 and $2.$ Define $\varrho_{n}(\Sigma):\mathcal{H}(2n)\to\mathbb{N}\cup\{+\infty\}$ by
@@ 3 The elliptic closed characteristics
$r\delta<\sum_{l=1,\,\frac{\theta_{l}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}}^{r}\left\{m_{j(s)}\frac{\theta_{j}}{\pi}\right\},$ (3.11)
@@
For any fix $a\in A(v)$ and $s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$, consider the inject map $p(N-s+1)=([(\tau_{j(s)},x_{j(s)})],2m_{j(s)})$ given by Theorem 2.4. Now, from Lemma 3.3, we can choose an $\hat{a}\in A(v)$ such that $\chi_{j(1)}(\hat{a})=1$. Then Theorem 3.6-$(1)$ implies $x_{j(1)}$ is an elliptic closed characteristic and
$\gamma_{j(1)}(\tau)$ $\simeq N_{1}(1,1)\diamond I_{2p_{0}}\diamond N_{1}(1,-1)^{\diamond p_{+}}\diamond N_{1}(-1,1)^{\diamond q_{-}}\diamond-I_{2q_{0}}$
with all $N_{2}(\lambda_{1},\nu_{1}),\ldots,N_{2}(\lambda_{r_{0}},\nu_{r_{0}})$ are rational normal form. Moreover, it is easy to see that $x_{j(1)}$ is irrationally elliptic when $\Sigma$ is non-degenerate. Since $\chi_{j(1)}(\hat{a})=1$, from Theorem 3.6-$(2)$, we know $\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$. Now, choose $\tilde{a}=-\hat{a}\in A(v)$, using Theorem 2.4 again, we still have $(\tilde{N},\tilde{m}_{1},\ldots,\tilde{m}_{q})$ , $\tilde{j}(s)$ and inject map $p(\tilde{N}-s+1)=([(\tau_{\tilde{j}(s)},x_{\tilde{j}(s)})],2\tilde{m}_{\tilde{j}(s)}),s\in\{1,\ldots,\varrho_{n}(\Sigma)\}$. Similarly, by applying Theorem 3.6-$(1)$ again, we obtain $x_{\tilde{j}(1)}$ has the same properties as $x_{j(1)}$. If $\tilde{j}(1)=j(1)$, from the definition of $\chi(a)$, we know that $\chi_{j(1)}(\hat{a})=1$ implies $\chi_{j(1)}(-\hat{a})=0$, hence $\chi_{\tilde{j}(1)}(\tilde{a})=\chi_{j(1)}(-\hat{a})=0$, but $\hat{i}(x_{\tilde{j}(1)},1)=\hat{i}(x_{j(1)},1)\in\mathbb{R}\setminus\mathbb{Q}$, this contradict with Theorem 3.6-$(2)$, hence this case will not happen, we always have $\tilde{j}(1)\neq j(1)$, then $x_{j(1)},x_{\tilde{j}(1)}$ are two different closed characteristics that possess the desired properties as required in Theorem 1.2. This completes the proof. ∎
@@
$\gamma_{j(2)}(\tau)$ $\simeq N_{1}(1,1)\diamond R(\theta_{1})\diamond\hat{M}$,
where $\hat{M}=R(\theta_{2}),N(\pm 1,a)$ with $a=\pm 1,0$. Since $\Sigma$ is non-degenerate, $\hat{M}$ must be $R(\theta_{2})$ with $\frac{\theta_{2}}{\pi}\in\mathbb{R}\setminus\mathbb{Q}$, then $x_{j(2)}$ must be irrational elliptic. In summary, we obtain that $x_{j(1)}$ is irrational elliptic and either $x_{j(2)}$ or $x_{j(3)}$ must be irrational elliptic. This completes the proof. ∎